# Recent advances in the numerical solution of multi-order fractional differential equations

### v1

@@ 6 The code fhbvm2 2
  • (127) is appropriate when the vector field is smooth everywhere;
@@
Concerning the evaluation of the abscissae and the weights for the quadrature(s):
@@ 7 Numerical tests
  • the code fhbvm2 2 here described, based on a FHBVM(22,22) method, which can handle the cases $\nu=1$ and $\nu=2$ in (1). In the case $\nu=1$, it will be used with the same parameters used for fhbvm2.
@@
$\mbox{mescd}~=~\max\left\{0,\,-\log_{10}\max_{n}\|(y(t_{n})-y_{n})./(1+|y(t_{n})|)\|_{\infty}\right\},$
@@
[…]
$y^{(0.5)}=\frac{1}{5}\left(\begin{array}[]{rr}-92&-87\\ -58&-63\end{array}\right)y-\frac{1}{10}\left(\begin{array}[]{r}67\\ 83\end{array}\right),\qquad t\in[0,100],\qquad y(0)=\left(\begin{array}[]{r}5\\ 10\end{array}\right).$ (130)
Besed on the choices in [12], the following parameters are used for the various codes, to construct the corresponding WPD:
  • fde12, fde12-10, fde_pi12_pc, fde_pi12_pc-10: ${\tt h\,}=10^{-i}$, $i=4,5,6$;
  • flmm2-1, flmm2-2, flmm2-3, fde_pi2_im: ${\tt h\,}=10^{-i}$, $i=2,3,4,5$;
  • fhbvm: ${\tt M\,}=5,\dots,10$;
  • fhbvm2, fhbvm2 2: ${\tt nu\,}=50$, ${\tt n\,}=1$, ${\tt N\,}=10i$, $i=5,\dots,10$.
The corresponding WPD is reported in Figure 2, from which one may conclude that:
  • fde12 can reach an accuracy of less than 4 mescd in about 700 sec. fde_pi12_pc  reaches a similar accuracy, but with a higher execution time (more than 1000 sec). For both codes, increasing the number of corrector iterations does not improve accuracy but only affects the execution time (in fact, more than 2000 sec are needed to achieve the same accuracy);
  • flmm2 can achieve about 5 mescd in about 1000 sec (the run with the fractional BDF2 is slightly less accurate), with a comparable execution time, whereas fde_pi2_im  can reach less than 3 mescd;
  • fhbvm, fhbvm2, and fhbvm2 2  do have a comparable performance (13-14 mescd in about 0.1 sec). In particular, as is expected, fhbvm2 and fhbvm2 2  do have practically the same behavior.
Figure 1: WPD for Problem (130).
Figure 2: WPD for Problem (131).
Problem 2
@@
Problem 3
The third problem is a multi-order variant of [12, Problem 7]:
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $-s(t,{\alpha}_{1},\beta)^{2}+y_{1}^{2}+g(t,{\alpha}_{2},\beta),\qquad t\in[0,2],$ (132)
$y_{1}(0)$ $=$ $y_{2}(0)~=~1,$
@@
Problem 4
This problem is a multi-order fractional version of the Brusselator problem, adapted from [27, Problem (32)]:
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $By_{1}-y_{1}^{2}y_{2},\qquad\qquad\qquad t\in[0,100],$ (133)
$y_{1}(0)$ $=$ $1.2,\qquad y_{2}(0)~=~2.8,$
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $a_{21}y_{1}y_{2}-a_{22}y_{2}^{2}-\frac{a_{23}y_{2}y_{3}}{1+\beta y_{2}}-r_{2}y_{2},\qquad\qquad t\in[0,T],$ (134)
$y_{3}^{({\alpha}_{2})}$ $=$ $a_{31}y_{1}y_{3}+\frac{a_{32}y_{2}y_{3}}{1+\beta y_{2}}-a_{33}y_{3}^{2}-r_{3}y_{3},\qquad\qquad y(0)~=~y^{0}\in\mathbb{R}^{3}.$

### v2

@@ 6.2 Convergence analysis
  • (152) is appropriate when the vector field is suitably regular everywhere;
@@
In the sequel we shall report a unified convergence analysis for a HBVM(k,s)(k,s) method assuming, without loss of generality, that  $T\leq 1$  in (1), in order to simplify the arguments. [^15] Differently, a time scaling can be done. By taking into account (5), (8), (12), (48), (58), and (63), and considering that, for all $n=1,\dots,N$,
$f_{i}(\sigma^{n}(ch_{n}))=\sum_{j\geq 0}P_{j}^{i}(c)\gamma_{ij}(\sigma^{n}),\qquad c\in[0,1],\qquad i=1,\dots,\nu,$
one obtains that the discrete solution exactly solves the sequence of perturbed problems
$\sigma_{in}^{({\alpha}_{i})}(ch_{n})$ $=$ $f_{i}(\sigma^{n}(ch_{n}))-\psi_{in}(ch_{n}),\qquad c\in[0,1],\qquad n=1,\dots,N,$
$\sigma_{i1}^{(\iota)}(0)$ $=$ $y_{i0}^{\iota}\in\mathbb{R}^{m_{i}},\qquad\iota=0,\dots,\ell-1,\qquad i=1,\dots,\nu,$
$\sigma^{n}$ $=$ $(\sigma_{1n}^{\top},\dots,\sigma_{\nu n}^{\top})^{\top}\in\mathbb{R}^{m},$ (155)
in place of (setting $y_{in}(ch_{n})=y_{i}(t_{n-1}+ch_{n})$,  $c\in[0,1]$):
$y_{in}^{({\alpha}_{i})}(ch_{n})$ $=$ $f_{i}(y^{n}(ch_{n})),\qquad c\in[0,1],\qquad n=1,\dots,N,$
$y_{i1}^{(\iota)}(0)$ $=$ $y_{i0}^{\iota}\in\mathbb{R}^{m_{i}},\qquad\iota=0,\dots,\ell-1,\qquad i=1,\dots,\nu,$
$y^{n}$ $=$ $(y_{1n}^{\top},\dots,y_{\nu n}^{\top})^{\top}\in\mathbb{R}^{m}.$ (156)
In (155), we have denoted
$\psi_{in}(ch_{n})~=~\sum_{j\geq s}P_{j}^{i}(c)\gamma_{ij}(\sigma^{n})\,+\,\sum_{j=0}^{s-1}P_{j}^{i}(c)\left[\gamma_{ij}(\sigma^{n})-\gamma_{ij}^{n}\right],$ (157)
the perturbation term, which is given by the sum of the truncation error in the expansion of $f(\sigma^{n}(ch_{n}))$ and of the quadrature error in the approximation of the Fourier coefficients $\gamma_{ij}(\sigma^{n})$ (see (12)). In this respect, hereafter we shall assume that $k$ is large enough so that the quadrature implemented in the FHBVM(k,s)(k,s) method has order at least $2s$. In such a case, following arguments similar to those in [12, Section 3], it can be shown that, under regularity assumptions on the functions $f_{i}$, one has that  [^16] Hereafter, $\|\cdot\|$ will denote the infinity norm.
$\Psi_{in}:=\max_{c\in[0,1]}\|\psi_{in}(ch_{n})\|,$ (158)
satisfies
  • for (151):
    $\Psi_{in}~\leq~\left\{\begin{array}[]{ccl}O(h_{1}^{{\alpha}_{i}}),&&\mbox{if}~n=1,\\[5.69054pt] O(h_{n}^{s}),&&\mbox{otherwise};\end{array}\right$.
  • for (152):             $\Psi_{in}~\leq~O(h^{s})$;
  • for (154):
    $\Psi_{in}~\leq~\left\{\begin{array}[]{ccl}O(h_{1}^{{\alpha}_{i}}),&&\mbox{if}~n=1,\\[5.69054pt] O(h_{n}^{s}),&&\mbox{if}~n=2,\dots,\mu,\\[5.69054pt] O(h^{s}),&&\mbox{otherwise.}\end{array}\right$.
Setting  $\Psi^{n}=\max_{i}\Psi_{in}$,
$e_{in}(ch_{n})=\|y_{in}(ch_{n})-\sigma_{in}(ch_{n})\|,\qquad e^{n}(ch_{n})=\max_{i}e_{in}(ch_{n})\qquad c\in[0,1],\qquad n=1,\dots,N,$
the piecewise functions
$e(t_{n-1}+ch_{n})\equiv e^{n}(ch_{n}),\qquad\Psi(t_{n-1}+ch_{n})\equiv\Psi^{n},\qquad c\in[0,1],\qquad n=1,\dots,N,$
${\alpha}=\min_{i}{\alpha}_{i}$, $L$ the maximum of the Lipschitz constants of the $f_{i}$, and considering that $\Gamma(x)^{-1}<1.2$, for $x>0$, one eventually arrives at the inequality:
$e(t)~\leq~1.2\int_{0}^{t}(t-\tau)^{{\alpha}-1}\left[Le(\tau)+\Psi(\tau)\right]\mathrm{d}\tau,\qquad t\in[0,T].$
We observe that, because of the choices (151)–(154) one deduces that:
 $\forall\,2,\dots,N~:~h_{n-1}\leq h_{n}.$
Consequently, we can assume that $\Psi(t)$ is non decreasing and therefore, by using the fractional version of the Gronwall Lemma [67, Corollary 2] (see also [66]), one deduces that:
$e(t)~\leq~K\int_{0}^{t}(t-\tau)^{{\alpha}-1}\Psi(\tau)\mathrm{d}\tau,$ (159)
with  $K=1.2\,E_{\alpha}(1.2\Gamma({\alpha})LT^{\alpha})$,  being  $E_{\alpha}(z)=\sum_{j\geq 0}\frac{z^{j}}{\Gamma({\alpha}j+1)}$  the one-parameter Mittag-Leffler function. Furthermore, considering that  [^17] Recall that  $h_{n}=t_{n}-t_{n-1}$  and, moreover,  $(t-t_{n-1})\leq T\leq 1$.
$\forall\,n=1,2,\dots,N,\mbox{\quad and\quad}t\geq t_{n}~:~\int_{t_{n-1}}^{t_{n}}(t-\tau)^{{\alpha}-1}\mathrm{d}\tau~\leq~\frac{h_{n}^{\tilde{\alpha}}}{{\tilde{\alpha}}}$,
with
${\tilde{\alpha}}=\min\{{\alpha},1\},$ (160)
and taking into account that $\Psi(t)$ is piecewise constant, one obtains:
$K\int_{0}^{t}(t-\tau)^{{\alpha}-1}\Psi(\tau)\mathrm{d}\tau~\leq~\frac{K}{{\tilde{\alpha}}}\sum_{k=1}^{N}h_{k}^{\tilde{\alpha}}\Psi^{k}.$
Moreover, by considering that for the graded mesh (151) and the graded part in (154)
$\sum_{k=1}^{n}h_{k}^{s+{\tilde{\alpha}}}=h_{n}^{s+{\tilde{\alpha}}}\sum_{k=0}^{n-1}r^{-k(s+{\tilde{\alpha}})}<\frac{h_{n}^{s+{\tilde{\alpha}}}}{1-r^{-s-{\tilde{\alpha}}}}$,
whereas for the uniform mesh $nh^{s+{\tilde{\alpha}}}\leq h^{s+{\tilde{\alpha}}-1}$, from (158) one eventually obtains that:
[…]
In Section 8 we report a few numerical tests, to assess (160)–(161) in the case (151).
Remark 8
From (161) one concludes that one can achieve arbitrarily high-order methods (though, suitably choosing the initial timestep, in the case of the graded or mixed meshes): as matter of fact, if $s\gg 1$, a spectrally accurate solution can be gained, as it happens in the ODE case [3].
7 The code fhbvm2 2
@@
[…]
Remark 9
We emphasize that the code fhbvm2 2  currently handles only the cases $\nu=1,2$, because the software in [41] only implements the case of two weighting functions. On the other hand, the extension of such software is not trivial, and is beyond the scope of this paper. Consequently, when an extended version of the code in [41] will be available, we shall adapt the code fhbvm2 2  accordingly, using the arguments explained in Sections 4 and 5.
@@ 8 Numerical tests
  • the code fhbvm2 2 here described, based on either a FHBVM(22,22) or a FHBVM(30,22) method, as explained above, which can handle the cases $\nu=1$ and $\nu=2$ in (1). In the case $\nu=1$, it will be used with the same parameters used for fhbvm2.
@@
$\mbox{mescd}~=~\max\left\{0,\,-\log_{10}\max_{n}\|(y(t_{n})-y_{n})./(1+|y(t_{n})|)\|\right\},$ (163)
@@
Remark 10
For the sake of completeness, we recall that in the definition (163) the absolute error $|y(t_{n})-y_{n}|$ is scaled, component-wise, by $(1+|y(t_{n})|)$, so that each entry of the vector $|y(t_{n})-y_{n}|./(1+|y(t_{n})|)$ is approximately the absolute error, if the corresponding entry of $|y(t_{n})|$ is much smaller than 1, or the relative error, if much larger that 1. This measure is widely used in the numerical comparison of numerical codes (see, e.g., the Test-Set for IVP Solvers [71]).
Figure 2: WPD for Problem (164).
@@
Problem 2
The next problem is a multi-order variant of [16, Problem 7]:
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $-s(t,{\alpha}_{1},\beta)^{2}+y_{1}^{2}+g(t,{\alpha}_{2},\beta),\qquad t\in[0,2],$
$y_{1}(0)$ $=$ $y_{2}(0)~=~1,$ (165)
@@
We use problem (165) also for assessing the convergence order (161), when using a graded mesh (151). For this purpose, we use a modification of the code fhbvm2 2  implementing a FHBVM(30,s)(30,s) method. We use the parameters ${\tt n}={\tt N}=N$, so that $h_{N}=2/N$, whereas the parameter nu is chosen so that $h_{N}^{s+{\alpha}}=h_{1}^{2{\alpha}}$. In fact, in such case (see (160)) ${\tilde{\alpha}}={\alpha}=0.2$, and the upper bound in (161) (graded case) turns out to be optimal. In doing this, we also consider that the parameter $r$ defining the graded mesh is given by (153), i.e., r=N/(N1)r=N/(N-1). As one may see, the numerical results listed in Table 3 are in good agreement with (161), since the convergence order appears to be approximately given by $s+0.2$, as expected, for the considered values $s=2,3,4$.
Table 3: Convergence results for HBVM(30,s)(30,s) used with a graded mesh with $h_{N}=2/N$.
$s=2$ $s=3$ $s=4$
NN error order error order error order
10 2.88e-02 4.36e-04 2.07e-06
20 6.89e-03 2.06 5.08e-05 3.10 1.17e-07 4.14
40 1.57e-03 2.13 5.68e-06 3.16 6.41e-09 4.19
80 3.48e-04 2.17 6.21e-07 3.19 3.43e-10 4.22
@@
Problem 3
This problem is a multi-order fractional version of the Brusselator problem, adapted from [34, Problem (32)]:
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $By_{1}-y_{1}^{2}y_{2},\qquad\qquad\qquad t\in[0,100],$
$y_{1}(0)$ $=$ $1.2,\qquad y_{2}(0)~=~2.8,$ (166)
@@
$y_{2}^{({\alpha}_{2})}$ $=$ $a_{21}y_{1}y_{2}-a_{22}y_{2}^{2}-\frac{a_{23}y_{2}y_{3}}{1+\beta y_{2}}-r_{2}y_{2},\qquad\qquad t\in[0,T],$
$y_{3}^{({\alpha}_{2})}$ $=$ $a_{31}y_{1}y_{3}+\frac{a_{32}y_{2}y_{3}}{1+\beta y_{2}}-a_{33}y_{3}^{2}-r_{3}y_{3},\qquad\qquad y(0)~=~y^{0}\in\mathbb{R}^{3}.$ (167)
@@
Figure 9: Numerical solution of Problem (167), $T=5000$ by using timestep $h=1$ (except the first 50 initial steps).
Remark 11
It is worth mentioning that the considered large width of the integration interval $(T=500)$ of problem (167) confirms the good stability properties of the FHBVM(30,22) method implemented in the code fhbvm2 2. However, to further confirm this fact, we consider $T=5000$ and, by using the coarsest mesh, among those used for the numerical tests $({\tt n\,}=1$, ${\tt nu\,}=50$, ${\tt N\,}=T)$, we obtain the plot in Figure 9, which still exhibits the correct periodic behavior.